A man divides $10,000 among three investments at 4%, 5% and 6% per annum respectively. His total income is $515 per year. His income from the 5% and 6% investments exceeds the income at 4% by $415. Find how much is invested at each rate.
Answer by mananth(16949)

A man divides $10,000 among three investments at 4%, 5% and 6% per annum respectively. His total income is $515 per year. His income from the 5% and 6% investments exceeds the income at 4% by $415. Find how much is invested at each rate.
..
let him invest $x at 4 %
income from 4% = 0.04x
...
investment at5%&6% = 10,000-x
..
income from 5%&6% = 0.04x+415
...
0.04x+415 = 515
0.04x=100
x=100/0.04
x= 10000/4
x= 2500
Balance investment = $7500 @ 5%&6%
income = 415
5/11*415 = 188.63 interest from 5%
amount invested = $3772.72
6/11 * 415= 226.36 interest from 6%
Amount invested = $3772.6
C H E C K
2500*0.04 + 3772*0.05 +3772*0.06= 515
..
m.ananth@hotmail.ca